# An Introduction to Zero-Order Optimization Techniques for Robotics

### v1

@@ 6       $x\leftarrow(1-\alpha\sum\limits_{k=1}^{K}w_{k})x+\alpha\sum\limits_{k=1}^{K}w_{k}x_{k}$
Another choice of weight could be to assign $w_{k}=\frac{1}{K_{e}}$ to the best $K_{e}$ samples and 00 to the others (by best, we mean samples with the lowest value $f(x_{k})$). As pointed out by [30], this elitist weighting provides an interesting connection to the well-known Cross-Entropy Method (CEM). We refer the reader to the appendix D for more details.
An important point that we overlooked is whether the covariance update rule in Equation (III-C) will maintain the matrix positive definite. If $10\leq\eta<$ 1, $w_{i}\geq 0$ and the weights sum to one, then the covariance matrix will always be positive definite provided that we started with a positive definite matrix [31]. However, in practice, there is no reason that the function $f$ would be such that the weights satisfy this property.
One possible solution is to design an update that operates on exponential coordinates so that the covariance matrix naturally stays positive definite. This is the idea behind xNES [32]. Thanks to the invariant properties of the natural gradient, both update rules align when the step size goes to zero.
Another approach could be to use the log-sum-exp transformation used for MPPI. Indeed, this transformation naturally maintains the weights positive and such that they sum to one. We omit the derivations as it is straightforward to show that applying the natural gradient to Equation (18) yields the update rule (III-C) with the Exponential Average weights (Eq (17)). We will refer to this algorithm as MPPI-CMA. Table I summarizes all the possible variations of the CMA algorithm depending on the choice of weight.
@@
MPPI-CMA Exponential Average (Eq 17)
@@
When dealing with TO, it is often preferred to design algorithms that can achieve linear complexity with respect to the time horizon $T$ (Equation (15)). This is typically achieved by exploiting the problem’s structure. One way to do so is to use a block-diagonal covariance matrix (with $T$ blocks of size $n_{u}$ where $n_{u}$ is the dimension of the control input), i.e. $\Sigma=\operatorname{blockdiag}(\Sigma_{1},\dots\Sigma_{T})$. Clearly, this makes the update rules of Predictive Sampling and MPPI easily implementable with linear complexity with respect to the time horizon.
However, this is less straightforward in the case of the CMA update. Specifically, Equation (III-C) will not necessarily maintain the block-diagonal structure of the covariance matrix. This is because $\Sigma$ is not the appropriate representation to work with. Instead of applying the natural gradient descent to the full matrix $\Sigma$, one can apply it to each block $\Sigma_{t}$. By construction, this approach will maintain the overall covariance matrix block-diagonal. Intuitively, the natural gradient update for each block matches Equation (III-C). We refer to this approach as block-diagonal CMA and leave a formal proof of this claim for future work. Importantly, this formulation allows the CMA update to be implemented in a sequential way, ensuring linear computational complexity with respect to the time horizon $T$.
@@ III-D Numerical experiments
We introduce four TO benchmark problems based on Hydrax [33]: Cartpole, DoubleCartPole, PushT and Humanoid. Control bounds are enforced with a penalty term in the cost function. In addition, for the PushT and Humanoid tasks, we rely on a zero-order spline interpolation to reduce the dimension of the problem. All results are averaged over six seeds. More details about the cost functions and additional visualizations are available online  [^1] https://github.com/ajordana/zoo-rob.
We compare the performance of Predictive Sampling, randomized smoothing, MPPI and MPPI-CMA. All algorithms use $2048$ samples per iteration. For MPPI and MPPI-CMA, we use a temperature of $\lambda=0.1$. For the CMA update, we use a full covariance as we did not observe that this was a computational bottleneck. Nevertheless, we provide results with block-diagonal CMA in Appendix E. Furthermore, we find that MPPI-CMA performs significantly better when using a separate step size for the mean update and the covariance update. For completeness, results using a shared step size are provided in Appendix E.
Figure 2 shows the cost evolution across iterations for each test problem. The performance of randomized smoothing varies across test problems and is highly sensitive to the step size choice. As expected, MPPI-CMA consistently outperforms MPPI. Interestingly, Predictive Sampling performs well given its simplicity.
@@
Figure 2: Cost according to the number of iterations for different test problems. The solid lines represent the median taken over six seeds.
@@ IV Policy optimization
One of the main limitations of TO is that it is not always straightforward to use it to derive a controller. Indeed, performing MPC implies solving TO problems online, which can be extremely challenging due to limited compute time. Even though recent works were able to achieve impressive zero-order MPC demonstrations, these performances have not yet matched those obtained by RL. The main strength of RL is that it relies solely on policy evaluation at runtime. All the computation is moved offline where the goal is to search for a policy that will maximize performance across all possible initial conditions. This can be written:
$F(\theta)=\mathbb{E}\left[J(\theta,s)\right]$ (29)
@@
$J(\theta,s)$ $=\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k})$ (30)
[…]
$\theta$ denote the parameters of the policy (e.g. the weights of a neural network). $r$ is the reward, $a$ is the control action. Note that we follow the RL formalism by trading the minimization of a cost function for the maximization of a reward. Furthermore, in contrast to TO, the optimization for each trajectory is done with an infinite horizon and a discount factor $\gamma$. We do not consider stochastic dynamics for simplicity as most robotic models are deterministic (e.g. locomotion or manipulation). However, the ideas presented in this section should extend to the stochastic dynamics case.
@@
Because the policy is deterministic, RL algorithms aiming to solve this formulation are called Deterministic Policy Gradient (DPG) algorithms [34]. While stochastic policies can be relevant in the context of games or partially observable settings, most robotic applications such as locomotion or manipulation can be solved with deterministic policies. In fact, in practice, even works that rely on RL algorithms derived with stochastic policies generally deploy a deterministic policy on the robot.
There exist two main sampling approaches to maximize (29). The first one is to directly rely on the zero-order techniques which implies sampling in the parameter space of a neural network, similarly to [35]. The second approach, used in RL, is to leverage the structure of the problem to sample in the action space. In this work, we are interested in the latter one. In the deterministic setting, the gradient can be estimated with the DPG theorem [34]:
$\partial_{\theta}J(\theta,s_{0})$ $=\sum_{k=0}^{\infty}\gamma^{k}\partial_{\theta}\pi(s_{k})\partial_{a}Q(s_{k},a_{k})$ (31)
@@
$Q(s,a)=\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k})$ (32)
[…]
@@
$\nabla_{\theta}F(\theta)$ $=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}\partial_{\theta}\pi_{\theta}(s_{k})\partial_{a}Q(s_{k},a_{k})\right]$ (33)
Note that we follow the formalism of [36] and express the policy gradient as an expectation over entire trajectories rather than making the $\gamma$-discounted state distribution explicit. Unfortunately, this formula cannot be directly implemented in an algorithm because of the presence of the $Q$-function. Indeed, the $Q$-function cannot be computed analytically. One solution is to estimate it with function approximation at each gradient descent step. Algorithms relying on this mechanism are called actor-critics. The actor refers to the policy while the critic refers to the $Q$-function. In practice, the $Q$-function is learned in a self supervised way via the Bellman equation. There exist many techniques to estimate the critic efficiently between gradient updates of the actor; we refer the reader to [11] for a comprehensive treatment of the subject.
@@
IV-A Gaussian smoothing on the $Q$-function
DDPG [38] originally proposed to learn the $Q$-function with a neural network and directly differentiate it to obtain $\partial_{a}Q(s_{k},a_{k})$. This idea led to state-of-the-art RL algorithms such as TD3 [39]. In light of the success of randomized smoothing, a natural idea could be to use Gaussian smoothing to estimate the gradient of the Q-function and use the following estimate for the gradient of the cost:
$\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}\partial_{\theta}\pi_{\theta}(s_{k})(Q(s_{k},a_{k}+\epsilon_{k})-Q(s_{k},a_{k}))\Sigma^{-1}\epsilon_{k}\right]$ (34)
where the expectation is taken with respect to both initial condition and the Gaussian random variables $\epsilon_{k}$. This provides us with a simple variation of actor-critics based on the DPG theorem. Algorithm (7) summarizes the procedure.
@@ 5       Update the critic, $Q$
We deliberately omit the rollout logic as well as the interplay between the actor update and the critic update. This is because our focus is solely on modifying the actor’s gradient update in DPG algorithms. We do not aim to provide a novel actor-critic logic. As we will see in the experimental section, this simple change can lead to competitive algorithms. An interesting variation of Algorithm (7) is to use the log-sum-exp transform and use exponential average weights in the actor update.
@@
One way to understand this algorithm is to interpret the partial derivative of the policy $\partial_{\theta}\pi_{\theta}(s)$ as a linear mapping from action space to the policy parameter space. Consequently, the algorithm can be understood as a way to map samples from action space into samples from the policy parameter space.
IV-B Numerical experiments
In this section, we investigate the benefits of estimating the gradient of the $Q$-function in DPG algorithms such as DDPG and TD3. Our implementation relies on CleanRL [40] and only modifies the actor’s update rule. Specifically, we investigate two smoothing approaches, the default randomized smoothing (RS) and the one relying on the log-sum-exp (LSE) transform. Consequently, we benchmark DDPG and TD3 against their smoothed counterparts: RS-DDPG, LSE-DDPG, RS-TD3, LSE-TD3. The performances are evaluated on seven Mujuco environments. For the sampling, we use 10 samples per update, with sampling noise following a Gaussian distribution with a standard deviation of $0.1$ (the covariance matrix is proportional to the identity matrix). Our implementation is available online  [^2] https://github.com/ajordana/zoo-rob.
Five runs are performed for each test problem. Figure 3 presents the normalized score across all runs. The episodic return for each environment is provided in Appendix F. The results indicate that both RS-DDPG and LSE-DDPG significantly outperform DDPG, which demonstrates the benefits of smoothing. For TD3, the improvements are not as clear. This could be explained by the fact that TD3 is already a very strong algorithm and that the margin of improvement on those benchmarks is limited. Additionally, our modifications were restricted to the actor’s update; achieving state-of-the-art performance would most likely require extensive hyperparameter tuning of the actor-critic mechanism. Nevertheless, the results show that deterministic policy gradient algorithms can benefit from smoothing.
@@ IV-B Numerical experiments
IV-C Connection to the Reinforce algorithm
One may ask how the update rule introduced in Equation (34) compares to the classical Reinforce algorithm. First, let’s recall how this algorithm is derived. In its most general form, Reinforce relies on the gradient of the following formulation.
$J(\theta,s_{0})=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{k}(S_{k},A_{k})\right]$ (35)
[…]
The policy is no longer a deterministic function but a parametrized distribution over the space of possible actions. Note that the expectation is now taken over both the initial conditions and random actions. We now use capitalized letter for actions and states to emphasize their stochastic nature. The stochastic policy gradient theorem states that
$\frac{\partial}{\partial\theta}J(\theta,s_{0})$ $=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}Q(S_{k},A_{k})\frac{\partial}{\partial\theta}\log(p_{\theta}(A_{k}|S_{k}))\right]$
Note that the definition of the $Q$-function is no longer the same as in the deterministic case. It is now defined as the expectation of stochastic rollouts starting at state $s$ with action $a$. In practice, in most RL algorithms, the actions are sampled from a Gaussian probability distribution with mean $\pi_{\theta}(s)$ and covariance $\Sigma$ where $\pi_{\theta}$ is a neural network. Hence, we have:
$a\sim\mathcal{N}(\pi_{\theta}(s),\Sigma)$ (36)

### v2

@@ 6 $x\leftarrow(1-\alpha\sum\limits_{k=1}^{K}w_{k})x+\alpha\sum\limits_{k=1}^{K}w_{k}x_{k}$
Another choice of weight could be to assign $w_{k}=\frac{1}{K_{e}}$ to the best $K_{e}$ samples and 0 to the others (by best, we mean samples with the lowest value $f(x_{k})$). As pointed out by [34], this elitist weighting provides an interesting connection to the well-known Cross-Entropy Method (CEM). We refer the reader to Appendix D for more details.
An important point that we overlooked is whether the covariance update rule in Equation (III-C) maintains the matrix positive definite. If $0\leq\eta<1$, $w_{i}\geq 0$, and the weights sum to one, then the covariance matrix is always positive definite, provided that we start with a positive definite matrix [35]. However, in practice, there is no reason that the function $f$ would be such that the weights satisfy this property.
One possible solution is to design an update that operates on exponential coordinates so that the covariance matrix naturally stays positive definite. This is the idea behind xNES [36]. Thanks to the invariant properties of the natural gradient, both update rules align when the step size approaches zero.
Another approach could be to use the log-sum-exp transformation used for MPPI. Indeed, this transformation naturally maintains the weights positive and ensures that they sum to one. We omit the derivations as it is straightforward to show that applying the natural gradient to Equation (20) yields the update rule from Equation (III-C) with the Exponential Average weights (Equation (19)). We refer to this algorithm as MPPI-CMA. Table I summarizes all the possible variations of the CMA algorithm, depending on the choice of weight.
@@
MPPI-CMA Exponential Average (Eq. 19)
@@
When dealing with TO, it is often preferred to design algorithms that can achieve linear complexity with respect to the time horizon $T$ (Equation (17)). This is typically achieved by exploiting the problem’s structure. One way to do so is to use a block-diagonal covariance matrix (with $T$ blocks of size $n_{u}$, where $n_{u}$ is the dimension of the control input), i.e. $\Sigma=\operatorname{blockdiag}(\Sigma_{1},\dots\Sigma_{T})$. Clearly, this makes the update rules of Predictive Sampling and MPPI easily implementable with linear complexity with respect to the time horizon.
However, this is less straightforward in the case of the CMA update. Specifically, Equation (III-C) does not necessarily maintain the block-diagonal structure of the covariance matrix. This is because $\Sigma$ is not the appropriate representation to work with. Instead of applying the natural gradient descent to the full matrix $\Sigma$, one can apply it to each block $\Sigma_{t}$. By construction, this approach maintains the overall covariance matrix block-diagonal. Intuitively, the natural gradient update for each block matches Equation (III-C). We refer to this approach as block-diagonal CMA and leave a formal proof of this claim for future work. Importantly, this formulation allows the CMA update to be implemented in a sequential way, ensuring linear computational complexity with respect to the time horizon $T$.
@@ III-D Numerical experiments
We introduce four TO benchmark problems based on Hydrax [37]: Cartpole, DoubleCartPole, PushT, and Humanoid. Control bounds are enforced with a penalty term in the cost function. All results are averaged over six seeds. More details about the cost functions and additional visualizations are available online  [^1] https://github.com/ajordana/zoo-rob.
We compare the performance of Predictive Sampling, randomized smoothing, MPPI, and MPPI-CMA. All algorithms use $2048$ samples per iteration. For MPPI and MPPI-CMA, we use a temperature of $\lambda=0.1$. For the CMA update, we use a full covariance as we did not observe that this was a computational bottleneck. Nevertheless, we provide results with block-diagonal CMA in Appendix E. Furthermore, we find that MPPI-CMA performs significantly better when using a separate step size for the mean update and the covariance update. For completeness, results using a shared step size are provided in Appendix E.
Figure 2 shows the cost evolution across iterations for each test problem. The performance of randomized smoothing varies across test problems and is highly sensitive to the choice of step size. As expected, MPPI-CMA consistently outperforms MPPI. Interestingly, Predictive Sampling performs well despite its simplicity.
@@
Figure 2: Cost according to the number of iterations for different TO test problems. The solid lines represent the median taken over six seeds.
@@ IV Policy optimization
One of the main limitations of TO is that it is not always straightforward to use it to derive a controller. Indeed, performing MPC implies solving TO problems online, which can be extremely challenging due to limited compute time. Even though recent works were able to achieve impressive zero-order MPC demonstrations [2, 3], these performances have not yet matched those obtained by RL [4, 5]. The main strength of current RL approaches is that they rely solely on policy evaluation at runtime. All the computations are moved offline, where the goal is to search for a policy that maximizes performance across all possible initial conditions. This can be written:
$F(\theta)=\mathbb{E}\left[J(\theta,s)\right],$ (31)
@@
$J(\theta,s)$ $=\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k}),$ (32)
$\mbox{s.t. }\,s_{0}=s,\,$ $s_{k+1}=f_{\text{dyn}}(s_{k},a_{k})\,\mbox{ and }\,a_{k}=\pi_{\theta}(s_{k}).$
$\theta$ denotes the parameters of the policy (e.g. the weights of a neural network). $r$ is the reward, $a$ is the control action. Note that we follow the RL formalism by trading the minimization of a cost function for the maximization of a reward. Furthermore, in contrast to TO, the optimization for each trajectory is done with an infinite horizon and a discount factor $\gamma$. We do not consider stochastic dynamics for simplicity, as most robotic models are deterministic (e.g. locomotion or manipulation). However, the ideas presented in this section should extend to the stochastic dynamics case. Because the policy is deterministic, RL algorithms aiming to solve formulation (31) are called Deterministic Policy Gradient (DPG) algorithms [38].
@@
Most RL algorithms rely on stochastic policies, therefore DPG algorithms are often introduced as a special case. While stochastic policies can be relevant in the context of games or partially observable settings, most robotic applications, such as locomotion or manipulation, can be solved with deterministic policies. In fact, in practice, many works that rely on RL algorithms derived with stochastic policies generally deploy a deterministic policy on the robot. Therefore, we chose to first introduce the deterministic case and then investigate how stochastic formulations can be understood as a random search aimed at solving the deterministic case.
There are two main sampling approaches to maximize (31). The first one is to directly rely on the zero-order techniques, which implies sampling in the parameter space of a neural network, similarly to [16, 17]. The second approach, used in RL, is to leverage the structure of the problem to sample from the action space. In this work, we are interested in the latter one.
IV-A Deterministic Policy Gradient
In the deterministic setting, the gradient can be estimated with the DPG theorem [38]:
$\partial_{\theta}J(\theta,s_{0})$ $=\sum_{k=0}^{\infty}\gamma^{k}\partial_{\theta}\pi(s_{k})\partial_{a}Q(s_{k},a_{k}),$ (33)
@@ IV-A Deterministic Policy Gradient
$Q(s,a)=\sum_{k=0}^{\infty}\gamma^{k}r(s_{k},a_{k}),$ (34)
$\mbox{s.t. }\,s_{0}=s,\;$ $a_{0}=a,\;s_{k+1}=f_{\text{dyn}}(s_{k},a_{k}),\;a_{k}=\pi_{\theta}(s_{k}).$
@@
$\hskip-5.69046pt\nabla_{\theta}F(\theta)$ $=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}\partial_{\theta}\pi_{\theta}(s_{k})\partial_{a}Q(s_{k},a_{k})\right].$ (35)
Note that we follow the formalism of [39] and express the policy gradient as an expectation over entire trajectories rather than making the $\gamma$-discounted state distribution explicit. Unfortunately, this formula cannot be directly implemented in an algorithm because of the presence of the $Q$-function. Indeed, the $Q$-function cannot be computed analytically. One solution is to estimate it with function approximation at each gradient descent step. Algorithms relying on this mechanism are called actor-critics. The actor refers to the policy, while the critic refers to the $Q$-function. In practice, the $Q$-function is learned in a self-supervised way via the Bellman equation. There exist many techniques to estimate the critic efficiently between gradient updates of the actor; we refer the reader to [11] for a comprehensive treatment of the subject.
@@
IV-B Gaussian smoothing on the $Q$-function
DDPG [41] originally proposed to learn the $Q$-function with a neural network and directly differentiate it to obtain $\partial_{a}Q(s_{k},a_{k})$. This idea led to state-of-the-art RL algorithms such as TD3 [42]. In light of the success of randomized smoothing, a natural idea is to use Gaussian smoothing to estimate the gradient of the Q-function and to use the following estimate for the gradient of the cost:
$\hskip-7.11317pt\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}\partial_{\theta}\pi_{\theta}(s_{k})(Q(s_{k},a_{k}+\epsilon_{k})-Q(s_{k},a_{k}))\Sigma^{-1}\epsilon_{k}\right],$ (36)
where the expectation is taken with respect to both the initial condition and the Gaussian random variables $\epsilon_{k}$. This provides us with a simple variation of actor-critics based on the DPG theorem. Algorithm (7) summarizes the procedure.
@@ 5   Update the critic, $Q$
We deliberately omit the rollout logic, as well as the interplay between the actor update and the critic update. This is because our focus is solely on modifying the actor’s gradient update in DPG algorithms. We do not aim to provide a novel actor-critic logic. As shown in the experimental section, this simple change can lead to competitive algorithms. An interesting variation of Algorithm 7 is to use the log-sum-exp transform and use exponential average weights in the actor update.
@@
One way to understand this algorithm is to interpret the partial derivative of the policy $\partial_{\theta}\pi_{\theta}(s)$ as a linear mapping from the action space to the policy parameter space. Consequently, the algorithm can be understood as a way to map samples from the action space into samples from the policy parameter space.
IV-C Numerical experiments
In this section, we investigate the benefits of estimating the gradient of the $Q$-function in DPG algorithms such as DDPG and TD3. Our implementation relies on CleanRL [43] and only modifies the actor’s update rule. Specifically, we investigate two smoothing approaches, the default randomized smoothing (RS) and the one relying on the log-sum-exp (LSE) transform. Consequently, we benchmark DDPG and TD3 against their smoothed counterparts: RS-DDPG, LSE-DDPG, RS-TD3, and LSE-TD3. The performances are evaluated in seven MuJoCo environments. For the sampling, we use 10 samples per update, with sampling noise following a Gaussian distribution with a standard deviation of $0.1$ (the covariance matrix is proportional to the identity matrix). Our implementation is available online  [^2] https://github.com/ajordana/zoo-rob.
Five runs are performed for each test problem. Figure 3 presents the normalized score across all runs. The episodic return for each environment is provided in Appendix F. The results indicate that both RS-DDPG and LSE-DDPG significantly outperform DDPG, which demonstrates the benefits of smoothing. For TD3, the improvements are not as clear. This could be explained by the fact that TD3 is already a very strong algorithm and that the margin for improvement on those benchmarks is limited. Additionally, our modifications were restricted to the actor’s update; achieving state-of-the-art performance would most likely require extensive hyperparameter tuning of the actor-critic mechanism. Nevertheless, the results show that deterministic policy gradient algorithms can benefit from smoothing.
@@ IV-C Numerical experiments
IV-D Connection to the Reinforce algorithm
One may ask how the update rule introduced in Equation (36) compares to the classical Reinforce algorithm [11]. First, let’s recall how this algorithm is derived. In its most general form, Reinforce relies on the gradient of the following formulation:
$J(\theta,s_{0})=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r_{k}(S_{k},A_{k})\right],$ (37)
$\mbox{s.t. }\;\;S_{0}$ $=s_{0}\;,A_{k}\sim p_{\theta}(\;.\;|S_{k}),\;S_{k+1}=f_{\text{dyn}}(S_{k},A_{k}).$
The policy is no longer a deterministic function but a parameterized distribution over the space of possible actions. Note that the expectation is taken over random actions. We now use capitalized letters for actions and states to emphasize their stochastic nature. The stochastic policy gradient theorem states that
$\frac{\partial}{\partial\theta}J(\theta,s_{0})$ $=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}Q(S_{k},A_{k})\frac{\partial}{\partial\theta}\log(p_{\theta}(A_{k}|S_{k}))\right].$
Note that the definition of the $Q$-function is no longer the same as in the deterministic case. It is now defined as the expectation of stochastic rollouts starting at state $s$ with action $a$. In practice, in most RL algorithms, the actions are sampled from a Gaussian probability distribution with mean $\pi_{\theta}(s)$ where $\pi_{\theta}$ is a neural network. While it is possible to also optimize the parameters of the covariance, we consider, for simplicity, a Gaussian distribution with a fixed covariance $\Sigma$. Hence, we have:
$a\sim\mathcal{N}(\pi_{\theta}(s),\Sigma).$ (38)